OpenStudy (anonymous):

Would someone be willing to explain a pre-calc problem involving solving trigonometric equations? Find all the solutions to the equation. 7 (sin^2)x - 14 sin x + 2 = -5

OpenStudy (phi):

this equation combines trig and algebra. for the moment , let y = sin(x) . This makes it easier to see: \[ 7 \sin^2x - 14 \sin x + 2 = -5 \\ 7 y^2 -14y +2 = -5\] First step: solve for y (you get two solutions). Do you know how ?

OpenStudy (anonymous):

Is it: 7y^2-14y+2=-5 7y^2-14y=-7 7(y^2-2y)=-7 7(y-1)(y-1) Solutions: y=1 and y=1?

OpenStudy (phi):

you have the correct factors, but your method is dubious. The standard way is to put it in standard form, like this 7y^2-14y=-7 7 y^2 -14y +7 = 0 (add 7 to both sides factor out 7: 7(y^2-2y+1) = 0 we can factor y^2-2y+1 into (y-1)(y-1) so 7 (y-1)(y-1) = 0 and that means y=1

OpenStudy (phi):

now remember y is short for sin x so y=1 means sin x = 1 now list all x's where sin x = 1

OpenStudy (anonymous):

y=1 at 90 degrees (pi/2). Is that what you mean by listing the x's?

OpenStudy (anonymous):

@phi ?

OpenStudy (phi):

yes, but if you want all the answers, remember the sine goes on forever. so every 2pi (or 360 degrees) later, it reaches 1 again. so the complete answer is \[ x= \frac{\pi}{2} + n2\pi \text{ for all integer n} \]